Vapour density of a metal chloride is 66. It`s oxide contains 53%of metal . The atomic mass of metal is

Grade:11

1 Answers

Arun
6 years ago

Equivalent of metal = Equivalent of chloride

53/n = 47/8 , E = 9.02 , Atomic mass of metal = n X 9.02

Molecular mass of chloride = 2 X 66 = n X 9.02 + n X 35.5

on solving n = 3

Atomic mass of metal = 9.02 X 3 = 27.06